The maximum magnetic field strength of an electromagnetic field is 5×10−6T. Calculate the maximum electric field strength if the wave is traveling in a medium in which the speed of the wave is 0.75c

1
Expert's answer
2022-01-23T16:14:00-0500

"E=\frac{B\varepsilon_0\mu_0}{\sqrt{1-\frac{v^2}{c^2}}}=7\cdot 10^6~\frac NC."